Which of the following bonds offers the highest current​ yield?

a. ​A(n) 7.32​%, ​20-year bond quoted at 100.885.

b. ​A(n)12.20​%, ​15-year bond quoted at 168.141.

c. ​A(n) 3.66%, ​18-year bond quoted at 50.442.
Answer #1

To compute the bonds current yield we shall be using the below formula:

= Annual interest / Market price of bond

Now we shall evaluate the above three options one by one:

a. Annual Interest = 100 x 7.32%

= 7.32

Market price of bond = 100.885

Now plugging these figures in the above mentioned formula we shall get:

= 7.32 / 100.885

= 7. 255786291 %

b. Annual Interest = 100 x 12.20%

= 12.20

Market price of bond = 168.141

Now plugging these figures in the above mentioned formula we shall get:

= 12.20 / 168.141

= 7. 25581506 %

c. Annual Interest = 100 x 3.66%

= 3.66

Market price of bond = 50.442

Now plugging these figures in the above mentioned formula we shall get:

= 3.66 / 50.442

= 7. 255858213 %

Now if we compare all the above three options we will find the highest current yield is obtained by option c.

Hence option c is the correct answer.

Feel free to ask in case of any query relating to this question